Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by umaa » Sun Feb 01, 2009 8:38 pm
A jar contains marbles of several different colors. Some marbles are to be taken from the jar one at a time, examined and the put back in the jar. What is the probability that exactly three of next five marbles taken from the jar will be blue?

(1) The probability that all of the next five marbles taken from the jar are blue is 243/3125

(2) The probability that none of the next five marbles taken from the jar are blue is 32/3125
Join the discussion
Source: — Problem Solving |

by sjd00d » Sun Feb 01, 2009 10:37 pm
should be D.

1. probably of next 5 marbles - 243/3125. probably p of getting blue in first try than p^5 = 243/3125, from this you can find value for p^3..Sufficient

2. probably of getting 5 marbles 1 - 32/3125, calculate from this as #1..Sufficient
Join the discussion

by ontopofit » Sun Feb 01, 2009 11:02 pm
yes IMO D.

1) tells there are 5 marbles in which 3 are blue;suff
2) tells the same thing that out of 5 2 are not blue;suff
Join the discussion

by aroon7 » Mon Feb 02, 2009 8:21 am
I think it is A

1) says: The probability that all of the next five marbles taken from the jar are blue is 243/3125

P(5blue) = 243/3125
p(blue)^5 = 243/3125

from this we can find out P(blue) and P(not blue)
ans will be P(blue)*P(blue)*P(blue)*P(not blue)*P(not blue)

Suff.

2) says 'The probability that none of the next five marbles taken from the jar are blue is 32/3125"
so this is 1-Prob(1 blue U 2 blue U 3 blue U 4 blue U 5blue) = 32/3125
we dont have any info about other marbles in the jar
so this is insuff.
Join the discussion

by dendude » Mon Feb 02, 2009 9:22 am
aroon7 wrote: 2) says 'The probability that none of the next five marbles taken from the jar are blue is 32/3125"
so this is 1-Prob(1 blue U 2 blue U 3 blue U 4 blue U 5blue) = 32/3125
we dont have any info about other marbles in the jar
so this is insuff.
This does not seem right.
From 32/3125 you can get [P(not Blue)]^5, from which you can get P(Blue), since 1-P(Blue) = P(not Blue)

So the Ans should be D
Join the discussion

by aroon7 » Mon Feb 02, 2009 10:18 am
dendude wrote:
aroon7 wrote: 2) says 'The probability that none of the next five marbles taken from the jar are blue is 32/3125"
so this is 1-Prob(1 blue U 2 blue U 3 blue U 4 blue U 5blue) = 32/3125
we dont have any info about other marbles in the jar
so this is insuff.
This does not seem right.
From 32/3125 you can get [P(not Blue)]^5, from which you can get P(Blue), since 1-P(Blue) = P(not Blue)

So the Ans should be D
dendue,

Following are the possible combinations:

b - blue
x - other color

BXXXX (5 times)
BBXXX (5C2 times)
BBBXX (5C3 times)
BBBBX (5c4 times)
BBBBB (1 time)

XXXXX (1 time)

so 32/3125 is prob of XXXXX
1 - 32/3125 gives you prob of other 5 combinations not just prob of BBBBB

if X is only one color we can calculate P(blue) but it is not explicitly stated...
so insuff
Join the discussion

by umaa » Mon Feb 02, 2009 10:29 am
The OA is D. But I'm not clear with the explanations you've given. It because I'm bad at probability. Can you please post a detailed explanation? Also, can you send me a link or a way to strengthen my probability area?
Join the discussion

by ontopofit » Mon Feb 02, 2009 10:59 am
yes aroon....ur right.
thats why 1-none will not give the probability of 5B's.

instead it is

(1-NB)^5.
hence D.
Join the discussion

by dendude » Mon Feb 02, 2009 12:29 pm
aroon7 wrote: dendue,

Following are the possible combinations:

b - blue
x - other color

BXXXX (5 times)
BBXXX (5C2 times)
BBBXX (5C3 times)
BBBBX (5c4 times)
BBBBB (1 time)

XXXXX (1 time)

so 32/3125 is prob of XXXXX
1 - 32/3125 gives you prob of other 5 combinations not just prob of BBBBB

if X is only one color we can calculate P(blue) but it is not explicitly stated...
so insuff
Aroon,
I dont think its necessary to worry about other colors.
According to me there are only two outcomes: Blue or Not Blue. We do not have to be concerned about how many other colors there are.

Stmt#2 gives,
P(NB)*P(NB)*P(NB)*P(NB)*P(NB) = 32/3125
i.e P(NB)^5 = 32/3125 from which you can derive P(NB) (which is 2/5)
P(B) can be easily derived from this and the value can be used to determine whats asked in the question.

So you can derive the required answer from either Stmts alone.
Hence D
Join the discussion

by krisraam » Mon Feb 02, 2009 12:35 pm
The Answer is D:

As we are replacing the balls after each try. The events are mutually exclusive.

Statement 1 says "The probability that all of the next five marbles taken from the jar are blue is 243/3125 "

Let the probabilty of taking a blue ball is x/y. If 5 you do that five times the probabilty is (x/y)^5 = 243/3125

So the probabilty of taking a blue ball is 3/5.

Probability of not taking a blue ball is 1-3/5 = 2/5.

Probabilty of taking a blue ball 3 times in 5 passes is (3/5)^3 * (2/5)^2

statement 1 is sufficient.

Statement 2 says "The probability that none of the next five marbles taken from the jar are blue is 32/3125"

Probabilty of not taking a blue ball in all of the 5 times is 32/3125.

Probability of not taking a blue ball is 2/5 ( this means 2 out of 5 times you will get a color other than blue)

Probabilty of taking a blue ball is 1-2/5 = 3/5

Statement 2 is also sufficient.

Statement 1 and 2 gives the same information.

Thanks
Raama
Join the discussion